OpenStudy (anonymous):

Mrs. Bird set-up her kindergarten classroom so that there was a right triangular area formed by the bookshelves If bookshelf A is 6 feet long and bookshelf B is 8 feet long, what is the length of bookshelf C?

OpenStudy (anonymous):

@mathman806 @Potatoes.ramu

OpenStudy (anonymous):

@oneinmillon

OpenStudy (anonymous):

do the same thing as last time a^2+b^2=c^2 (6*6)+(8*8)=(c*c) 36+64=c^2 100=c^2 10=c

OpenStudy (anonymous):

If R = 25 cm and S = 7 cm, what is the length of T?

OpenStudy (anonymous):

is R the longest side?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

a^2+7^2=25^2 a^2+49=625 a^=576 a=24

OpenStudy (anonymous):

if x = 17 cm and y = 8 cm what is the length of z??

OpenStudy (anonymous):

what is the longest side?

OpenStudy (anonymous):

X

OpenStudy (anonymous):

8^2+b^2=17^2 64+b^2=289 b^2=225 b=15